Find the probability of answering the first 3 questions on a test correctly if random guesses are made, assume that all the questions are multiple choice, each with 5 possible answers.
Answer #1

n = 3

p = 1/5 = 0.2

It is a binomial distribution.

P(X = x) = nCx * px * (1 - p)n - x

p(X = 3) = 3C3 * (0.2)^3 * (0.8)^0 = 0.008